OpenStudy (anonymous):

5^y=k Find 5^(y+1) , in terms of k Medal given

OpenStudy (anonymous):

we have : 5^y=k 5^y+1 K+1

OpenStudy (anonymous):

no that is not the answer

OpenStudy (anonymous):

\[5^y=k\\ 5^{y+1}=5\times 5^y\] is a start

OpenStudy (anonymous):

How to proceed

OpenStudy (anonymous):

since \(5^y=k\) then \(5\times 5^y=5k\)

OpenStudy (anonymous):

and so \(5^{y+1}=5k\)

OpenStudy (anonymous):

i mean second

OpenStudy (anonymous):

In your second step why did u multiply 5^y with 5

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

because \(\large 5\times 5^y=5^{y+1}\)

OpenStudy (anonymous):

got it thanks :)